OpenStudy (anonymous):

Help Please? Complete the following ordered pairs for the equation. 4x + 5y = -22 (_, -2) (2, _) (7, _)

OpenStudy (anonymous):

kellie what happen

OpenStudy (anonymous):

4x + 5y = -22 (_,-2).....y = -2 so sub -2 in for y to find x 4x + 5(-2) = -22 4x - 10 = -22 4x = 22 + 10 4x = 32 x = 8 so the first one is (8,-2) understand ?

OpenStudy (anonymous):

lala...I can't just stay on one question all day...I float around. If you have to ask me more, just tag me and I will be there when I can. Right now I am with bluedog

OpenStudy (anonymous):

Okay that makes sense sort of. how would i set up the next one because this time i believe the 2 is x

OpenStudy (anonymous):

but you were helping me first and at my question its the last one

OpenStudy (anonymous):

hold on lala...let me finish with bluedog

OpenStudy (anonymous):

uhh ok

OpenStudy (anonymous):

(2,_) x = 2 so sub 2 in for x 4x + 5y = -22 4(2) + 5y = -22 8 + 5y = -22 5y = -22 - 8 5y = - 30 y = -6 so the second one would be (2,-6) you gettin it now ?

OpenStudy (anonymous):

Okay so what im getting at is i would take for the next one 4(7)+5=-22 then 28+5=-22 is that right so far?

OpenStudy (anonymous):

is -10 the right answer @kelliegirl33 ?

OpenStudy (anonymous):

x = 7 4x + 5y = -22 4(7) + 5y = -22 --- don't forget to put your y 28 + 5y = -22 5y = -22 - 28 5y = - 50 y = -10 yep...you got it. (7,-10) good job :)